Five accounts are randomly sampled without replacement from a box of forty accounts. The

box contains thirty accounts that have a debit balance. What is the probability of selecting at

least one account with a credit balance?


1
Expert's answer
2022-05-12T18:26:53-0400

Let "X=" the number of accounts with a credit balance.


"P(X\ge 1)=1-P(X=0)"

"=1-\dfrac{\dbinom{40-10}{0}\dbinom{30}{5}}{\dbinom{40}{5}}"

"=1-\dfrac{1(142506)}{658008}\approx0.783428"